OpenStudy (kimberly_pr):

Evaluate the summation of 2 n plus 5, from n equals 1 to 12 Will medal and/or fan!

OpenStudy (kimberly_pr):

Answer choices: 29 36 216 432

OpenStudy (kimberly_pr):

@Mehek14

OpenStudy (kimberly_pr):

@Qwertty123

OpenStudy (kimberly_pr):

@ganeshie8 @pooja195

OpenStudy (kimberly_pr):

@Atsie

OpenStudy (dumbcow):

\[\sum_{n=1}^{12} 2n+5 = 2 \sum_{n=1}^{12}n + \sum_{n=1}^{12}5 = 2 (\frac{12(12+1)}{2}) + 12(5) = 216\]

OpenStudy (kimberly_pr):

Thank you so much!! :) So you basically just have to sum across? I don't really get how to solve this

OpenStudy (dumbcow):

right you break it up into smaller summations you just need the formula: \[\sum_{i=1}^{N} i = \frac{N(N+1)}{2}\]

OpenStudy (kimberly_pr):

Oh okay, I'll write that equation down. I'm in virtual school and they don't teach these concepts well :/